OpenStudy (anonymous):

Please help! Will fan and medal! Your friend brings hot chocolate to school in the cylindrical thermos and drinks from her cylindrical mug shown.

OpenStudy (anonymous):

1. Your friend fills the thermos with hot chocolate to a height of 10 inches. How much does she bring to school? Leave your answer in terms of π. Volume equation for a cylinder is pi * r^2 * h. r=2in, h = 10in. π * 2in * 2in * 10in = π40in. She brings π40in^3. 2. The inside bottom of the mug is 4 inches from the top. Find the volume of the mug. Leave your answer in terms of π. Volume equation for a cylinder is pi * r^2 * h. r = 1.5in, h = 4in. π * 1.5in * 1.5in * 4in = π * 2.25in * 4in = π9in^3. 3. Each time your friend fills the mug, she fills it 80% full and then drinks it completely before filling up again. How much hot chocolate does she pour into the mug each time? Leave your answer in terms of π. 80% of 9 in = 7.2in. 80% = 0.80. 0.8 times 9in = 7.2in. Each time that your friend fills the mug up, she fills it up π7.2in^3. 4. Is the mug more than half full after the last pour from the thermos? Justify your answer.

OpenStudy (anonymous):

How do I find how many pours the thermos has?